To test the significance of regression you need to evaluate the variance (s^2). What is the variance?

Single choice.





1
Expert's answer
2021-07-26T15:27:35-0400

Sample variance ("s^2") is a measure of the degree to which the numbers in a list are spread out. If the numbers in a list are all close to the expected values, the variance will be small. If they are far away, the variance will be large. Sample variance is given by the equation


"s^2=\dfrac{\sum (x_i-\bar x)^2}{n-1}"

where n is the number of categories.

and "\bar x =\dfrac{\sum _{i=1}^nx_i}{n}"